OpenStudy (anonymous):

Elimination: y=6x^2 7x+y=3

OpenStudy (anonymous):

Im certain this one isnt as simple. At least I hope so...

OpenStudy (solomonzelman):

Rearrange the second equation \(\color{#000000}{ \displaystyle 7x+y=3 }\) \(\color{#000000}{ \displaystyle y=-7x+3 }\) substitute this into the first equation \(\color{#000000}{ \displaystyle y=6x^2 }\) \(\color{#000000}{ \displaystyle -7x+3=6x^2 }\) \(\color{#000000}{ \displaystyle 0=6x^2+7x-3 }\) Solve this quadratic equation for x.

OpenStudy (solomonzelman):

Then, after you solve for x, you can solve for y.

OpenStudy (anonymous):

oh, didnt realize that I could substitue the 6x2 for y

OpenStudy (anonymous):

well thanks again

OpenStudy (solomonzelman):

np